A+A+A+A=b6
C+C=A
single digit number which fits both equations

Answer by Edwin McCravy(20054)


A+A+A+A=b6
C+C=A

I don't get the "b6"

Maybe you mean that the "b" is a digit also.

4+4+4+4=16
2+2=4

????

Edwin